We have looked a little more at the SRM situation. We aligned the SRM, and then aligned the oplev, so that we had a convenient monitor of the optic's motion.
When we use the _COMM channels, which are the usual ones on the IFO_ALIGN screen, the pitch slider makes pitch motion, but the yaw slider makes the oplev spot move ~45degrees from horizontal.
However, when we use the bias channels that are in the front end model, parallel to the ASC path, pitch moves pitch, and yaw moves pure yaw.
So, we conclude that the SRM coils are fine, and there is something funny going on with the slow part of the actuation.
Koji restarted the slow computer susaux, and burt restored it, but that did not fix the situation. We went inside and looked at all of the ribbon cable connections, and pushed them all in, but that also has not fixed things.
We have been looking at D010001-b, the coil driver board, and we think that's where the summing resistor network between the slow bias slider, and the coil outputs from the fast model exists. (It's not 100% clear, but we're confident that that's what is going on).
Tomorrow, we will pull the SRM's coil driver board, and see if any of the components in the slow slider path, before the summing point, look burned / broken / bad.